OpenStudy (anonymous):

What is the area of a rectangle with a length of feet and a width of feet? A. square feet B. square feet C. square feet D. square feet

OpenStudy (anonymous):

http://prntscr.com/9mlgo8

OpenStudy (anonymous):

@civicsiscool44

OpenStudy (anonymous):

@ChantySquirrel1129**

OpenStudy (anonymous):

@Study312

OpenStudy (civicsiscool44):

sorry i am confused

OpenStudy (anonymous):

@jabez177

OpenStudy (civicsiscool44):

@zepdrix

OpenStudy (anonymous):

@lolacole12

OpenStudy (study312):

I ain't good at area with fractions, sry.

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

oh it a trick question it doesnt matter what you answer there all the same

OpenStudy (anonymous):

It's basically asking you length times width.

OpenStudy (anonymous):

To find the answer you do need to multiply length times width. But, how do you do that? One way is to make an improper fraction out of each and then multiply them. 6 &2/5 = 6*5+2= 32/5 do you see how that worked? Now do 4& 1/6 then multiply them to find your answer.